# Black holes and black regions, horizons and barriers in Lorentzian manifolds

### v1

@@ Abstract.
We prove that if $\mathscr{S}$ is a time-oriented null hypersurface of a Lorentzian $n$-manifold $(M,g)$, the causal world-lines, which intersect transversally $\mathscr{S}$ and are time-oriented in a compatible way, cross the hypersurface all in the same direction, the other being forbidden. Even if it is known that a smooth event horizon (in the sense of Penrose, Hawking and Ellis) is a null hypersurface and has the above semi-permeability property, at the best of our knowledge, it was not known so far that the latter is a mere consequence of the former. Our result leads to the concepts of barriers (= null hypersurfaces separating the space-time into disjoint regions) and black regions (= time-oriented regions bounded by barriers). These objects naturally include (smooth) event horizons and (smoothly bounded) black holes. Since barriers are defined by two simple properties – the merely local property of “nullity” combined with the global property of “separating the space-time” – we expect they may be used to simplify computations for locating static and/or dynamic horizons in numerical computations.
@@ 1. Introduction
$g_{S}=\left(1-\frac{2m}{r}\right)du^{2}-2du\vee dr-r^{2}d\xi^{2}-r^{2}\sin^{2}\xi d\psi^{2}\ .$ (1.1)
@@ 3.1. The semi-permeability of the dressed barriers
$g_{x_{o}}(E^{\perp}_{x_{o}},E^{\perp}_{x_{o}})=0\ ,\qquad g_{x_{o}}(\mathbb{T}_{x_{o}},E^{\perp}_{x_{o}})=-1\ .$ (3.1)
@@
$\begin{split}&g(\mathbb{T}_{y},\mathbb{T}_{y})w^{1}>-\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)+\\ &\hskip 71.13188pt+\sqrt{\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)^{2}+g(\mathbb{T}_{y},\mathbb{T}_{y})\sum_{j=2}^{n-1}(w^{j})^{2}}\ ,\\ &\hskip 170.71652pt\text{or}\end{split}$ (3.3)
$\begin{split}&g(\mathbb{T}_{y},\mathbb{T}_{y})w^{1}<-\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)-\\ &\hskip 71.13188pt-\sqrt{\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)^{2}+g(\mathbb{T}_{y},\mathbb{T}_{y})\sum_{j=2}^{n-1}(w^{j})^{2}}\ .\end{split}$ (3.4)
@@
Under the hypotheses of Theorem 3.2, there is no causal world-line $\mathcal{C}\subset\mathscr{U}$, which is $\mathscr{T}{}$-time oriented and crosses $\mathscr{S}$ at $x_{o}$ in the direction of $E^{\perp}]$.
@@ 3.2. The semi-permeability property of time-oriented null hypersurfaces
$\mathbb{T}=\widetilde{E}-v\widetilde{E}^{\perp}\$. (3.5)
@@
$g_{x}(E^{\perp},E)=-1\qquad\text{for any}\ x\in\mathscr{S}\ .$ (3.6)
Then no oriented causal world-line $\mathcal{C}\subset M$ satisfies all of the following three conditions:
  • (a)
    it intersects transversally $\mathscr{S}$ at the point $x_{o}=\mathcal{C}\cap\mathscr{S}$;
  • (b)
    the velocity $v=\dot{\gamma}(s_{o})$ at $x_{o}$ of a parameterisation $\gamma$ of its orientation $\gamma]$ is in ${\mathscr{T}{}}^{(\mathscr{S})}_{x_{o}}$;
  • (c)
    it crosses $\mathscr{S}$ at $x_{o}$ in the direction of $E^{\perp}]$.
Proof. Assume that there is an oriented causal world-line $\mathcal{C}$ satisfying (a) – (c) and consider a neighbourhood $\mathscr{U}$ of $x_{o}$ and a vector field $\mathbb{T}$ on $\mathscr{U}$ as in Lemma 3.4. Denote also by $\mathscr{T}{}$ the time-orientation of $\mathscr{U}_{+}$ determined by $\mathbb{T}|_{\mathscr{U}_{+}}$. Then the oriented arc $\mathcal{C}^{\prime}=\mathcal{C}\cap\mathscr{U}$ is an oriented causal world-line for the Lorentzian manifold $(\mathscr{U},g|_{\mathscr{U}})$. From (b) and the fact that $\mathbb{T}|_{\mathscr{U}\cap\mathscr{S}}=E|_{\mathscr{U}\cap\mathscr{S}}$, the orientation of $\mathcal{C}^{\prime}$ gives a $\mathscr{T}{}$-time orientation on $\mathcal{C}^{\prime}\cap\mathscr{U}_{+}$, while by (a) and (c) $\mathcal{C}^{\prime}$ crosses $\mathscr{S}\cap\mathscr{U}$ in the direction of $E^{\perp}]$. This contradicts Corollary 3.3.
@@ 4. A few (classical and not) examples of black holes and barriers
    $g^{-1}(dF,dF)|_{\{F=R\}}>0\qquad\bigg(\ \text{resp.}\ <0\ ,\ \ =0\ \bigg)\ .$ (4.1)
@@ 4.1. The inner and outer barriers of Kerr-Newman black holes
$g_{KN}:=-\frac{1}{2}(\rho^{2}+a^{2}\cos^{2}\!\xi)\left(d\xi^{2}+\sin^{2}\!\xi d\psi^{2}\right)-\\ -\left(dv-a\sin^{2}\!\xi d\psi\right)\vee\left(d\rho-a\sin^{2}\!\xi d\psi\right)+\frac{1}{2}\left(1-\frac{2m\rho-e^{2}}{\rho^{2}+a^{2}\cos^{2}\!\xi}\right)\big(dv-a\sin^{2}\!\xi d\psi\big)^{2}\ ,$ (4.2)
@@
$E_{(+)}:=-\frac{\partial}{\partial\psi}+\frac{2}{a}(m+\sqrt{m^{2}-a^{2}-Q^{2}})\frac{\partial}{\partial v}\bigg|_{\mathscr{S}_{R_{+}}}\ ,\\ E_{(-)}:=-\frac{\partial}{\partial\psi}+\frac{2}{a}(m-\sqrt{m^{2}-a^{2}-Q^{2}})\frac{\partial}{\partial v}\bigg|_{\mathscr{S}_{R_{-}}}\ ,$ (4.3)
@@ 4.2. Barriers and horizons of Myers and Perry’s n-dimensional black holes
$g_{MP}=-\frac{1}{2}(\rho^{2}+a^{2}\cos^{2}\!\xi)(dxi^{2}+\sin^{2}\!\xi\,d\psi^{2})+\rho^{2}\cos^{2}\!\xi\,g_{S^{n-4}}-\\ -(dv-a\sin^{2}\!\xi\,d\psi)\vee(d\rho-a\sin^{2}\!\xi\,d\psi)+\frac{1}{2}\left(1-\frac{m}{\rho^{n-5}(\rho^{2}+a^{2}\cos^{2}\!\xi)}\right)(dv-a\sin^{2}\!\xi d\psi)^{2}\ ,$ (4.4)
@@
$0=g_{MP}^{-1}(dF,dF)\big|_{\{F=0\}}=G(\theta_{i},\xi,\psi,R)(R^{n-5}(R^{2}+a^{2})-m)\\ \text{where}\ G(\theta_{i},\xi,\psi,R)\ \text{is an appropriate nowhere vanishing function on}\ M\ .$ (4.5)
@@ 4.3. Barriers in Ricci flat manifolds with Kerr type optical structures
$\Delta\varphi+\kappa\frac{8\varphi}{\left(1+\kappa(x^{2}+y^{2})\right)^{2}}=0\ ,$ (4.6)
@@
$g^{(\kappa,\varphi)}:= -2\kappa\frac{\rho^{2}+\varphi^{2}}{(1+\kappa(x^{2}+y^{2}))^{2}}(dx^{2}+dy^{2})+\\ +\big(dv-\frac{\partial\varphi}{\partial y}dx+\frac{\partial\varphi}{\partial x}dy\big)\vee\bigg(dv+d\rho+\frac{1}{2}\left(-1+\frac{2m\,\rho}{\rho^{2}+\varphi^{2}}\right)\big(dv-\frac{\partial\varphi}{\partial y}dx+\frac{\partial\varphi}{\partial x}dy\big)\bigg)\ ,$ (4.7)
@@
$\begin{split}&\varphi(r,\psi)=-\left(a_{0}\frac{\varphi_{0}(r)}{\varphi_{0}(r_{o})}+\sum^{\infty}_{n=1}\frac{\varphi_{n}(r)}{\varphi_{n}(r_{o})}\big(a_{n}\cos(n\psi)+b_{n}\sin(n\psi)\big)\right)\ ,\\ &\text{with}\ \ \ \varphi_{n}(r):=\left(1-\frac{2r^{2}}{(1+r^{2})}\frac{1}{n+1}\right)r^{n}\ \ \bigg(\text{resp. }\ \ \ \varphi_{n}(r):=\left(\frac{1+r^{2}}{1-r^{2}}+n\right)r^{n}\ \ \bigg)\ .\end{split}$ (4.8)
@@
$g^{(\kappa,\varphi)-1}\bigg(\frac{\partial R}{\partial x}dx+\frac{\partial R}{\partial y}dy+\frac{\partial R}{\partial v}dv,\frac{\partial R}{\partial x}dx+\frac{\partial R}{\partial y}dy+\frac{\partial R}{\partial v}dv\bigg)\bigg|_{\rho=R(x,y,v)}=\\ =g^{(\kappa,\varphi)-1}(d\rho,d\rho)\big|_{\rho=R(x,y,v)}\ .$ (4.9)

### v2

@@ Abstract.
We prove that if $\mathscr{S}$ is a time-oriented null hypersurface of a Lorentzian $n$-manifold $(M,g)$, the causal world-lines, which intersect transversally $\mathscr{S}$ and are time-oriented in a compatible way, cross the hypersurface all in the same direction, the other being forbidden. Even if it is known that a smooth event horizon (in the sense of Penrose, Hawking and Ellis) is a null hypersurface and has the above semi-permeability property, at the best of our knowledge, in the literature it was not stated so far that the latter is a mere consequence of the former. Our result leads to the concepts of barriers (= null hypersurfaces separating the space-time into disjoint regions) and black regions (= time-oriented regions bounded by barriers). These objects naturally include (smooth) event horizons and (smoothly bounded) black holes. Since barriers are defined by two simple properties – the merely local property of “nullity” combined with the global property of “separating the space-time” – we expect they may be used to simplify computations for locating static and/or dynamic horizons in numerical computations.
@@ 1. Introduction
(1.1) $g_{S}=\left(1-\frac{2m}{r}\right)du^{2}-2du\vee dr-r^{2}d\xi^{2}-r^{2}\sin^{2}\xi d\psi^{2}\ .$
@@ 3.1. The semi-permeability of the dressed barriers
(3.1) $g_{x_{o}}(E^{\perp}_{x_{o}},E^{\perp}_{x_{o}})=0\ ,\qquad g_{x_{o}}(\mathbb{T}_{x_{o}},E^{\perp}_{x_{o}})=-1\ .$
@@
(3.3) $\begin{split}&g(\mathbb{T}_{y},\mathbb{T}_{y})w^{1}>-\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)+\\ &\hskip 71.13188pt+\sqrt{\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)^{2}+g(\mathbb{T}_{y},\mathbb{T}_{y})\sum_{j=2}^{n-1}(w^{j})^{2}}\ ,\\ &\hskip 170.71652pt\text{or}\end{split}$
(3.4) $\begin{split}&g(\mathbb{T}_{y},\mathbb{T}_{y})w^{1}<-\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)-\\ &\hskip 71.13188pt-\sqrt{\left(g_{y}(\mathbb{T}_{y},f_{0}(y))w^{0}+\sum_{k=2}^{n-1}g_{y}(\mathbb{T}_{y},f_{k}(y))w^{k}\right)^{2}+g(\mathbb{T}_{y},\mathbb{T}_{y})\sum_{j=2}^{n-1}(w^{j})^{2}}\ .\end{split}$
@@
Under the hypotheses of Lemma 3.2, there is no causal world-line $\mathcal{C}\subset\mathscr{U}$, which is $\mathscr{T}{}$-time oriented and crosses $\mathscr{S}$ at $x_{o}$ in the direction of $E^{\perp}]$.
@@ 3.2. The semi-permeability property of time-oriented null hypersurfaces
(3.5) $\mathbb{T}=\widetilde{E}-v\widetilde{E}^{\perp}\$.
@@
(3.6) $g_{x}(E^{\perp},E)=-1\qquad\text{for any}\ x\in\mathscr{S}\ .$
Given $x_{o}\in\mathscr{S}$ and a neighbourhood $\mathscr{U}\subset M$ of $x_{o}$, equipped with a vector field $\mathbb{T}$ as in Lemma 3.4 and with the time-orientation $\mathscr{T}$ on $\mathscr{U}_{+}$ determined by $\mathbb{T}$, no oriented causal world-line $\mathcal{C}\subset M$ crosses $\mathscr{S}$ at $x_{o}$ in the direction of $E^{\perp}]$ (resp. $E^{\perp}]$) if its orientation is compatible with the $\mathscr{T}$-time-orientation (resp. $(-\mathscr{T})$-time-orientation) in $\mathscr{U}_{+}$.
Proof. If the orientation is compatible with the $\mathscr{T}$-time-orientation, the arc $\mathcal{C}^{\prime}=\mathcal{C}\cap\mathscr{U}$ is a $\mathscr{T}$-time-oriented causal world-line for the Lorentzian manifold $(\mathscr{U},g|_{\mathscr{U}})$ and cannot cross $\mathscr{S}\cap\mathscr{U}$ in the direction of $E^{\perp}]$ by Corollary 3.3. An analogous conclusion occurs when the orientation is compatible with the $(-\mathscr{T})$-time-orientation.
@@ 4. A few (classical and not) examples of black holes and barriers
    (4.1) $g^{-1}(dF,dF)|_{\{F=R\}}>0\qquad\bigg(\ \text{resp.}\ <0\ ,\ \ =0\ \bigg)\ .$
@@ 4.1. The inner and outer barriers of Kerr-Newman black holes
(4.2) $g_{KN}:=-\frac{1}{2}(\rho^{2}+a^{2}\cos^{2}\!\xi)\left(d\xi^{2}+\sin^{2}\!\xi d\psi^{2}\right)-\\ -\left(dv-a\sin^{2}\!\xi d\psi\right)\vee\left(d\rho-a\sin^{2}\!\xi d\psi\right)+\frac{1}{2}\left(1-\frac{2m\rho-e^{2}}{\rho^{2}+a^{2}\cos^{2}\!\xi}\right)\big(dv-a\sin^{2}\!\xi d\psi\big)^{2}\ ,$
@@
(4.3) $E_{(+)}:=-\frac{\partial}{\partial\psi}+\frac{2}{a}(m+\sqrt{m^{2}-a^{2}-Q^{2}})\frac{\partial}{\partial v}\bigg|_{\mathscr{S}_{R_{+}}}\ ,\\ E_{(-)}:=-\frac{\partial}{\partial\psi}+\frac{2}{a}(m-\sqrt{m^{2}-a^{2}-Q^{2}})\frac{\partial}{\partial v}\bigg|_{\mathscr{S}_{R_{-}}}\ ,$
@@ 4.2. Barriers and horizons of Myers and Perry’s n-dimensional black holes
(4.4) $g_{MP}=-\frac{1}{2}(\rho^{2}+a^{2}\cos^{2}\!\xi)(dxi^{2}+\sin^{2}\!\xi\,d\psi^{2})+\rho^{2}\cos^{2}\!\xi\,g_{S^{n-4}}-\\ -(dv-a\sin^{2}\!\xi\,d\psi)\vee(d\rho-a\sin^{2}\!\xi\,d\psi)+\frac{1}{2}\left(1-\frac{m}{\rho^{n-5}(\rho^{2}+a^{2}\cos^{2}\!\xi)}\right)(dv-a\sin^{2}\!\xi d\psi)^{2}\ ,$
@@
(4.5) $0=g_{MP}^{-1}(dF,dF)\big|_{\{F=0\}}=G(\theta_{i},\xi,\psi,R)(R^{n-5}(R^{2}+a^{2})-m)\\ \text{where}\ G(\theta_{i},\xi,\psi,R)\ \text{is an appropriate nowhere vanishing function on}\ M\ .$
@@ 4.3. Barriers in Ricci flat manifolds with Kerr type optical structures
(4.6) $\Delta\varphi+\kappa\frac{8\varphi}{\left(1+\kappa(x^{2}+y^{2})\right)^{2}}=0\ ,$
@@
(4.7) $g^{(\kappa,\varphi)}:= -2\kappa\frac{\rho^{2}+\varphi^{2}}{(1+\kappa(x^{2}+y^{2}))^{2}}(dx^{2}+dy^{2})+\\ +\big(dv-\frac{\partial\varphi}{\partial y}dx+\frac{\partial\varphi}{\partial x}dy\big)\vee\bigg(dv+d\rho+\frac{1}{2}\left(-1+\frac{2m\,\rho}{\rho^{2}+\varphi^{2}}\right)\big(dv-\frac{\partial\varphi}{\partial y}dx+\frac{\partial\varphi}{\partial x}dy\big)\bigg)\ ,$
@@
(4.8) $\begin{split}&\varphi(r,\psi)=-\left(a_{0}\frac{\varphi_{0}(r)}{\varphi_{0}(r_{o})}+\sum^{\infty}_{n=1}\frac{\varphi_{n}(r)}{\varphi_{n}(r_{o})}\big(a_{n}\cos(n\psi)+b_{n}\sin(n\psi)\big)\right)\ ,\\ &\text{with}\ \ \ \varphi_{n}(r):=\left(1-\frac{2r^{2}}{(1+r^{2})}\frac{1}{n+1}\right)r^{n}\ \ \bigg(\text{resp. }\ \ \ \varphi_{n}(r):=\left(\frac{1+r^{2}}{1-r^{2}}+n\right)r^{n}\ \ \bigg)\ .\end{split}$
@@
(4.9) $g^{(\kappa,\varphi)-1}\bigg(\frac{\partial R}{\partial x}dx+\frac{\partial R}{\partial y}dy+\frac{\partial R}{\partial v}dv,\frac{\partial R}{\partial x}dx+\frac{\partial R}{\partial y}dy+\frac{\partial R}{\partial v}dv\bigg)\bigg|_{\rho=R(x,y,v)}=\\ =g^{(\kappa,\varphi)-1}(d\rho,d\rho)\big|_{\rho=R(x,y,v)}\ .$